Construction of the Fourfold Cover of the Mathieu Group M_{22}

Ibrahim A. I. Suleiman Robert A. Wilson

Abstract

We give an explicit construction for the two faithful, irreducible, 16dimensional representations of 4^{extstyle.}M_{22} over the field mathrmGF (49). Then we extend them to the 32-dimensional representation of 4^{extstyle.}M_{22}:2 over mathrmGF (7). Explicit matrices are given on page ??.